# Generative Autoregressive Transformers for Model-Agnostic Federated MRI Reconstruction

### v1

@@ Abstract
Although learning-based models hold great promise for MRI reconstruction, single-site models built on limited local datasets often suffer from poor generalization. This challenge has spurred interest in collaborative model training on multi-site datasets via federated learning (FL)– a privacy-preserving framework that aggregates model updates instead of sharing imaging data. Conventional FL builds a global model by aggregating locally trained model weights, inherently constraining all sites to a homogeneous model architecture. This rigid homogeneity requirement forces sites to forgo architectures tailored to their compute infrastructure and application-specific demands. Consequently, existing FL methods for MRI reconstruction fail to support model-heterogeneous settings, where individual sites are allowed to use distinct architectures. To overcome this fundamental limitation, here we introduce FedGAT, a novel model-agnostic FL technique based on generative autoregressive transformers. FedGAT decentralizes the training of a global generative prior that captures the distribution of multi-site MR images. For enhanced fidelity, we propose a novel site-prompted GAT prior that controllably synthesizes MR images from desired sites via autoregressive prediction across spatial scales. Each site then trains its site-specific reconstruction model–using its preferred architecture–on a hybrid dataset comprising the local MRI dataset and GAT-generated synthetic MRI datasets for other sites. Comprehensive experiments on multi-institutional datasets demonstrate that FedGAT supports flexible collaborations while enjoying superior within-site and across-site reconstruction performance compared to state-of-the-art FL baselines.
@@
MRI, reconstruction, federated learning, model agnostic, generative, autoregressive, transformer.
@@ 1 Introduction
Magnetic Resonance Imaging (MRI) is an indispensable diagnostic modality due to its exceptional soft-tissue contrast. Yet, characteristically long scan times in MRI exams can disrupt clinical workflows and reduce patient comfort [1]. A mainstream approach to improve scan efficiency relies on reconstruction of MR images from undersampled acquisitions [2, 3]. Learning-based reconstruction models have shown particular promise in recovering high-quality images from undersampled k-space data [4, 5, 6, 7, 8, 9, 10, 11, 12]. Despite their potential, models trained on local datasets available at individual sites often struggle to represent rare tissue features and fail to generalize effectively across sites [13, 14, 15, 16, 17, 18, 19, 20, 21]. As such, the development of robust models remains hindered by the limited availability of diverse training datasets in solitary institutions due to economic costs and privacy concerns [22, 23, 24]. This limitation underscores the pressing need for collaborative approaches to build models on multi-site MRI datasets while preserving patient privacy [25, 26].
@@
Federated learning (FL) is an emergent paradigm for collaborative model training in privacy-sensitive domains such as medical imaging [27, 28]. By exchanging model weights as opposed to raw imaging data, FL enables knowledge transfer among sites while alleviating potential privacy risks [29]. Recognizing the transformative potential of FL, several recent studies on MRI reconstruction have proposed FL methods for building models on multi-site datasets [30, 31, 32, 33, 34, 35]. A common theme among these studies has been their adherence to the conventional FL framework based on a server-client topology [29]. In this framework, a globally shared model is selected via agreement across sites to serve as the foundation for both cross-site knowledge transfer and eventual execution of the reconstruction task. Over multiple communication rounds, local model copies trained on local datasets available at individual sites are aggregated on the server to update the global model [36]. The global model is eventually expected to possess knowledge on the distribution of multi-site MRI datasets and thereby improve generalization over single-site models trained exclusively on local datasets [30].
@@
The conventional FL framework promises enhanced generalization in MRI reconstruction, while imposing a strict requirement for all participating sites to adopt a homogeneous model architecture as the basis of model aggregation [36]. However, imaging sites often differ substantially in terms of their computational resources and the complexity of their reconstruction tasks as influenced by variations in distribution of local datasets. These differences frequently lead sites to favor distinct model architectures, ranging from convolutional networks that can be preferred to attain high local precision under relatively low compute budgets [30, 31], transformers that can be preferred to boost sensitivity to long-range contextual features when high compute budgets are available [37, 38], to physics-driven unrolled networks that can be preferred to boost model reliability when training sets are relatively scarce albeit moderate-to-high compute budgets are available [34, 33]. Consequently, the architectural constraints in conventional FL refrain sites from leveraging models optimized for their specific requirements. This apparent lack of architectural freedom has important practical implications including discouraging resource-constrained sites from participating in collaborations, compromising performance at sites with more advanced resources due to use of relatively simple models, and ultimately stifling innovations in MRI reconstruction models. Therefore, enabling model-agnostic FL—where sites can collaborate without adhering to a homogeneous model architecture—remains a critical open challenge [39].
@@
In this study, we introduce a novel FL technique based on generative autoregressive transformers (FedGAT), which to our knowledge is the first model-agnostic FL approach in the literature for MRI reconstruction. Unlike conventional FL methods, FedGAT supports model-heterogeneous settings by decoupling the process of cross-site knowledge transfer on data distributions, from the process of building reconstruction models (Fig. 1). It mediates knowledge transfer via a global generative prior designed to synthesize multi-site MR images. To avoid the limitations of existing adversarial priors (i.e., training instability and poor quality) and diffusion priors (i.e., suboptimal convergence and heavy compute burden), here we propose a novel GAT prior that autoregressively synthesizes MR images by predicting feature maps across spatial scales, guided by a site prompt to preserve site-specific characteristics. Afterwards, site-specific reconstruction models are locally trained on a hybrid dataset, combining local MRI datasets available at each site with synthetic datasets from remaining sites generated by the GAT prior. Comprehensive demonstrations on multi-institutional MRI datasets indicate that FedGAT functions effectively in model-heterogeneous settings, overcoming a significant challenge in enabling privacy-preserving collaborations across diverse institutions. Code for FedGAT is available at https://github.com/icon-lab/FedGAT.
@@
Figure 1: FedGAT devises a two-tier strategy to collaboratively train heterogeneous models for MRI reconstruction. (a) The first tier conducts decentralized training of a global prior $\text{GAT}_{\theta_{\text{GAT}}}$ that captures the distribution of multi-site MR images. The global prior is built on a novel generative autoregressive transformer that formulates image generation as an autoregressive process across growing spatial scales, and that employs a site prompt sp for controllable image synthesis from desired sites. (b) The second tier conducts local training of site-specific reconstruction models $H^{k}_{{\phi}^{k}}$ ($k$: site index). To achieve an optimal balance between generalization and site-specific performance, training at each site is performed on a hybrid dataset comprising the local dataset and synthetic datasets from other sites generated via the GAT prior.
@@ 2.1 Learning-Based MRI Reconstruction
Learning-based models promise improved performance and reliability over traditional methods in MRI reconstruction [5]. Yet, realizing this promise requires copious amounts of training data including diverse samples, since learning-based models poorly represent rare features and pathologies in their training sets [40]. Unfortunately, curating broad training sets is unfeasible in many applications given the high costs of scanning and subject recruitment [41]. These costs bear the common practice of model training on limited datasets available locally in solitary imaging sites, and such single-site models can inevitably suffer from suboptimal generalization [42]. Numerous strategies have been devised over the years to alleviate the unwanted influences of dataset scarcity, including transfer learning [7], unpaired learning [15, 43], semi-supervised learning [44], and self-supervised learning [45, 46]. However, these canonical strategies rely on centralized model training following aggregation of datasets in a central repository, which incurs significant privacy concerns and maintenance costs [25].
@@ 2.2 Federated Learning for MRI Reconstruction
Federated learning (FL) has emerged as a privacy-preserving collaboration framework that distributes the costs of model training across healthcare institutions [25]. Abandoning explicit sharing of sensitive imaging data, FL promotes cross-site knowledge transfer via exchange of model weights instead. In conventional FL, this exchange is attained by first training copies of a global model on the local datasets available at individual sites, and then aggregating the local copies into the shared global model on a server [27]. Several recent MRI studies have successfully adopted this conventional FL framework to boost generalization performance of reconstruction models [30, 31, 32]. Further improvements in site-specific performance have also been sought by employing personalized FL strategies to cope with differences in the distribution of MRI datasets across sites, such as partial model aggregation [32], test-time adaptation [31], and feature map normalization [42].
@@
To enable collaborative training of heterogeneous models across sites, here we introduce FedGAT as the first model-agnostic FL method for MRI reconstruction. Unlike conventional FL that transfers knowledge by communicating weights of the reconstruction model, FedGAT leverages a unique approach that decouples the process of knowledge transfer across sites from the process of building site-specific reconstruction models, and that mediates knowledge transfer via a global generative prior for multi-site MR images. Note that previous FL methods in machine learning that build generative priors have commonly proposed adversarial priors that can suffer from poor training stability and image quality [49, 31], and diffusion priors that can suffer from suboptimal convergence that can lead to residual image noise and prolonged run times [50]. In contrast, here we introduce a novel prior based on generative autoregressive transformers that efficiently synthesizes MR images via autoregressive prediction of feature maps across growing spatial scales, under guidance from a site prompt to preserve site-specific characteristics in MR images. At each site, a site-specific reconstruction model is then locally trained on a hybrid dataset comprising both the local dataset and synthetic datasets from other sites generated via the global prior. These technical advances enable FedGAT to operate seamlessly in model-heterogeneous settings, overcoming a critical barrier towards privacy-preserving collaborations across diverse institutions.
@@ 3.1 Conventional FL for MRI Reconstruction
$\widehat{\mathbf{x}}^{k}=H_{\phi^{*}}(\mathcal{F}^{-1}\mathbf{y}^{k},\mathcal{A}^{k}),$ (5)
where $\widehat{\mathbf{x}}^{k}$ is the reconstructed image, and $\mathbf{y}^{k}$ are the respective undersampled k-space data.
@@ 3.2 Federated Generative Autoregressive Transformers
FedGAT is a model-agnostic FL technique that enables collaborative training of heterogeneous model architectures across sites by leveraging a decoupled two-tier strategy. During the first tier, using a server-client topology, a site-prompted global prior based on generative autoregressive transformers (GAT) is decentrally trained so as to capture the distribution of multi-site MR images (Fig. 1a). The FL server sends the trained global prior to individual sites at the end of the first tier. During the second tier, site-specific reconstruction models are locally trained with the aid of this global GAT prior (Fig. 1b). At each site, a reconstruction model of preferred architecture is pre-trained on the local dataset, and later fine-tuned on a hybrid dataset containing both the local dataset and synthetic datasets from other sites generated via the prior. The architecture of the site-prompted GAT prior, and image synthesis and two-tier training procedures for FedGAT are detailed in subsections below.
@@ 3.2.1 Architecture of the Site-Prompted GAT Prior
Figure 2: Architecture of the proposed site-prompted GAT prior. (a) The GAT prior embodies a variational autoencoder (VAE), whose encoder module maps an input MR image onto a set of discrete token maps $\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{S}$ across $S$ spatial scales, and whose decoder module recovers the input MR image back from the derived token maps. (b) The GAT prior also embodies a transformer module that builds an autoregressive prior on the multi-scale discrete token maps captured by the VAE. This module expresses generation of multi-scale discrete token maps as an autoregressive process across spatial scales, wherein a higher-scale token map $\mathbf{f}_{s}$ is predicted given the collection of token maps at earlier scales $\mathbf{f}_{<s}$:=assign\,:=\,:=$\{\mathbf{f}_{1},...,\mathbf{f}_{s-1}\}$. To preserve site-specific attributes in synthetic MR images, a site-token $\mathbf{st}$ derived from a site prompt is included in the autoregressive transformer module.
@@
Autoregressive Transformer Module: While the VAE encoder extracts discrete token maps residually across spatial scales, it does not explicitly examine the statistical relationships among these multi-scale maps [56]. To obtain a correlate for the joint distribution of $\mathbf{f}_{1},..,\mathbf{f}_{S}$, here we propose to build an autoregressive prior on multi-scale discrete token maps via the autoregressive transformer module. In particular, the transformer module aims to predict the sequence of tokens $\mathbf{f}_{s}$ at the $s$th scale, given as input the sequence of tokens aggregated across lower scales $\mathbf{f}_{<s}$:=assign\,:=\,:=$\{\mathbf{f}_{1},...,\mathbf{f}_{s-1}\}\in[V]^{T_{s-1}}$ ($T_{s-1}=\sum_{o=1}^{s-1}h_{o}w_{o}$ denoting the number of tokens at lower scales), and a learnable site-prompt $\mathbf{sp}(k)\in\mathbb{R}^{1\times d}$ used to initialize a site token, $\mathbf{st}$=\,=\,=$\mathbf{sp}(k)$. To maintain spatial correspondence between the token maps at consecutive scales, the token maps in $\mathbf{f}_{<s}$ are individually embedded after upsampling to the dimensions of the subsequent scale:
@@
where $\mathbf{e}_{o}\in\mathbb{R}^{h_{o+1}w_{o+1}\times d}$ and $\mathbf{e}_{<s}$:=assign\,:=\,:=$\{\mathbf{e}_{1},...,\mathbf{e}_{s-1}\}\in\mathbb{R}^{(T_{s}-1)\times d}$. Afterwards, the site token is pooled with a learnable projection of these embedded token maps, and additively combined with learnable position and scale encodings:
@@
To predict $\mathbf{f}_{s}$, the hidden representation $\mathbf{h}_{0}$ is projected through $L$ transformer blocks, each containing multi-head self-attention (MHSA) and multi-later perceptron (MLP) layers interleaved with normalization and residual connections [55]. Yet, unlike normalization layers in conventional transformers, here we employ adaptive layer normalization (AdaLN) in order to induce site-specific processing of hidden representations [61]. In particular, we propose to modulate the statistics of $\mathbf{h}$ via learnable functions of the site token $\mathbf{st}$:

### v2

@@ Abstract
Although learning-based models hold great promise for MRI reconstruction, single-site models built on limited local datasets often suffer from poor generalization. This challenge has spurred interest in collaborative model training on multi-site datasets via federated learning (FL)—a privacy-preserving framework that aggregates model updates instead of sharing imaging data. Conventional FL aggregates locally trained model weights into a global model, inherently constraining all sites to use a homogeneous model architecture. This rigidity forces sites to compromise on architectures tailored to their compute resources and application-specific needs, making conventional FL unsuitable for model-heterogeneous settings where each site may prefer a distinct architecture. To overcome this limitation, we introduce FedGAT, a novel model-agnostic FL technique based on generative autoregressive transformers. FedGAT decentralizes the training of a global generative prior that learns the distribution of multi-site MR images. For high-fidelity synthesis, we propose a novel site-prompted GAT prior that controllably synthesizes realistic MR images from desired sites via autoregressive prediction across spatial scales. Each site then trains its own reconstruction model–using an architecture of its choice–on a hybrid dataset augmenting its local MRI dataset with GAT-generated synthetic MR images emulating datasets from other sites. This hybrid training strategy enables site-specific reconstruction models to generalize more effectively across diverse data distributions while preserving data privacy. Comprehensive experiments on multi-institutional datasets demonstrate that FedGAT enables flexible, model-heterogeneous collaborations and achieves superior within-site and cross-site reconstruction performance compared to state-of-the-art FL baselines.
@@
MRI reconstruction, federated learning, generative, autoregressive transformer.
@@ 1 Introduction
Magnetic Resonance Imaging (MRI) is an indispensable diagnostic modality due to its exceptional soft-tissue contrast. Yet, characteristically long scan times in MRI exams can disrupt clinical workflows and reduce patient comfort [1]. A mainstream approach to improve scan efficiency relies on reconstruction of MR images from undersampled acquisitions [2]. Learning-based reconstruction models have shown particular promise in recovering high-quality images from undersampled k-space data [3, 4, 5, 6, 7, 8, 9, 10, 11]. Despite their potential, models trained on local datasets available at individual sites often struggle to represent rare tissue features and fail to generalize effectively across sites [12, 13, 14, 15, 16, 17, 18, 19, 20]. As such, the development of robust models remains hindered by the limited availability of diverse training datasets in solitary institutions due to economic costs and privacy concerns [21, 22, 23]. This limitation underscores the pressing need for collaborative approaches to build models on multi-site MRI datasets while preserving patient privacy [24, 25].
@@
Federated learning (FL) is an emergent paradigm for collaborative model training in privacy-sensitive domains such as medical imaging [26, 27]. By exchanging model weights as opposed to raw imaging data, FL enables knowledge transfer among sites while alleviating potential privacy risks [28, 29]. Recognizing the transformative potential of FL, several recent studies on MRI reconstruction have proposed FL methods for building models on multi-site datasets [30, 31, 32, 33, 34, 35]. A common theme among these studies has been their adherence to the conventional FL framework based on a server-client topology [36]. In this framework, a globally shared model is selected via agreement across sites to serve as the foundation for both cross-site knowledge transfer and eventual execution of the reconstruction task. Over multiple communication rounds, local model copies trained on local datasets available at individual sites are aggregated on the server to update the global model [37]. The global model is eventually expected to possess knowledge on the distribution of multi-site MRI datasets and thereby improve generalization over single-site models trained exclusively on local datasets [30].
@@
The conventional FL framework promises enhanced generalization in MRI reconstruction, while imposing a strict requirement for all participating sites to adopt a homogeneous model architecture as the basis of model aggregation [37]. However, imaging sites often differ substantially in terms of their computational resources and the characteristics of their reconstruction tasks as influenced by variations in the distribution of local datasets. These differences frequently lead sites to favor distinct model architectures, ranging from convolutional networks that can be preferred to attain high local precision under relatively low compute budgets [30, 31], transformers that can be preferred to boost sensitivity to long-range contextual features when high compute budgets are available [38, 39], to physics-driven unrolled networks that can be preferred to boost model reliability when training sets are relatively scarce albeit moderate-to-high compute budgets are available [34, 33]. Consequently, the architectural constraints in conventional FL refrain sites from leveraging models optimized for their specific needs. This apparent lack of architectural freedom has important practical implications including discouraging resource-constrained sites from participating in collaborations, compromising performance at sites with more advanced resources due to use of relatively simple models, and ultimately stifling innovations in MRI reconstruction models. Therefore, enabling model-agnostic FL—where sites can collaborate without adhering to a homogeneous model architecture—remains a critical open challenge [40].
@@
In this study, we introduce a novel FL technique based on generative autoregressive transformers (FedGAT), which to our knowledge is the first model-agnostic FL approach in the literature for MRI reconstruction. Unlike conventional FL methods that require consistent model architectures across sites, FedGAT supports model-heterogeneous settings via a two-tier strategy that decouples cross-site knowledge transfer, from building of reconstruction models (Fig. 1). To mediate knowledge transfer without sharing private data, in the first tier, FedGAT trains a global generative prior —a generative autoregressive transformer (GAT)— capable of synthesizing realistic multi-site MR images. Although adversarial and diffusive methods have been common in medical image synthesis [41, 42, 43], adversarial priors are amenable to training instabilities [44], diffusion priors are amenable to high computational complexity [45], potentially compromising convergence, image fidelity and practical utility in FL settings [31, 46]. To address these limitations, we propose a novel GAT prior that autoregressively predicts feature maps at multiple spatial scales, conditioned on site-specific prompts to preserve the unique characteristics of each site’s data. The proposed prior balances efficiency and fidelity while capturing diverse data distributions across sites within a unified model. Having a trained copy of the GAT prior, in the second tier, each site builds its own reconstruction model using a hybrid dataset: augmenting its own local MRI dataset with synthetic images generated by the GAT prior to emulate datasets from other sites. This hybrid training ensures that site-specific reconstruction models are exposed to broader data variability while preserving data privacy. Comprehensive experiments on multi-institutional MRI datasets demonstrate that FedGAT enables effective federated learning in model-heterogeneous settings, representing an important advance toward privacy-preserving collaborations across diverse institutions. Code for FedGAT is available at https://github.com/icon-lab/FedGAT.
@@
Figure 1: FedGAT implements a two-tier strategy to collaboratively train heterogeneous models for MRI reconstruction across multiple imaging sites. (a) The first tier conducts decentralized training of a global prior $\text{GAT}_{\theta_{\text{GAT}}}$ that learns the distribution of multi-site MR images. This global prior is based on a novel generative autoregressive transformer that formulates image generation as an autoregressive process across growing spatial scales, guided by a site prompt sp to preserve site-specific characteristics. (b) The second tier performs local training of site-specific reconstruction models $H^{k}_{{\phi}^{k}}$ ($k$: site index). Each model is trained on a hybrid dataset that combines the local MRI dataset with GAT-generated synthetic images representing datasets from other sites, enabling broader generalization while maintaining data privacy.
@@ 2.1 Learning-Based MRI Reconstruction
Learning-based models promise improved performance and reliability over traditional methods in MRI reconstruction [4]. Yet, realizing this promise requires copious amounts of training data including diverse samples, since learning-based models poorly represent rare features and pathologies in their training sets [47, 48, 49]. Unfortunately, curating broad training sets is unfeasible in many applications given the high costs of scanning and subject recruitment [50]. These costs bear the common practice of model training on limited datasets available locally in solitary imaging sites, and such single-site models can inevitably suffer from suboptimal generalization [51]. Numerous strategies have been devised over the years to alleviate the unwanted influences of dataset scarcity, including transfer learning [6], unpaired learning [14, 52], semi-supervised learning [53], and self-supervised learning [54, 55]. However, these canonical strategies rely on centralized model training following aggregation of datasets in a central repository, which incurs significant privacy concerns and maintenance costs [24].
@@ 2.2 Federated Learning for MRI Reconstruction
Federated learning (FL) has emerged as a privacy-preserving collaboration framework that distributes the costs of model training across healthcare institutions [24]. Abandoning explicit sharing of sensitive imaging data, FL promotes cross-site knowledge transfer via exchange of model weights instead. In conventional FL, this exchange is attained by first training copies of a global model on the local datasets available at individual sites, and then aggregating the local copies into the shared global model on a server [26]. Several recent MRI studies have successfully adopted this conventional FL framework to boost generalization performance of reconstruction models [30, 31, 32]. Further improvements in site-specific performance have also been sought by employing personalized FL strategies to cope with differences in the distribution of MRI datasets across sites, including partial model aggregation [32], test-time adaptation [31], and feature map normalization [51].
@@
To enable collaborative training of heterogeneous models across sites, here we introduce FedGAT as the first model-agnostic FL method for MRI reconstruction. Unlike conventional FL that transfers knowledge by communicating weights of the reconstruction model, FedGAT leverages a unique approach that decouples the process of knowledge transfer across sites from the process of building site-specific reconstruction models, and that mediates knowledge transfer via a global generative prior for multi-site MR images. Note that previous FL methods in machine learning that build generative priors have commonly proposed adversarial priors that can suffer from poor training stability and image quality [58, 31], and diffusion priors that can suffer from suboptimal convergence that can lead to residual image noise and prolonged run times [59]. In contrast, here we introduce a novel prior based on generative autoregressive transformers that efficiently synthesizes MR images via autoregressive prediction of feature maps across growing spatial scales, under guidance from a site prompt to preserve site-specific characteristics in MR images. At each site, a site-specific reconstruction model is then locally trained on a hybrid dataset comprising both the local dataset and synthetic datasets from other sites generated via the global prior. These technical advances enable FedGAT to operate effectively in model-heterogeneous settings, overcoming a critical barrier towards collaborations across diverse institutions.
@@ 3.1 Conventional FL for MRI Reconstruction
$\widehat{\mathbf{x}}\hskip 1.00006pt^{k}=H_{\phi^{*}}(\mathcal{F}^{-1}\mathbf{y}^{k},\mathcal{A}^{k}),$ (5)
where $\widehat{\mathbf{x}}\hskip 1.00006pt^{k}$ is the reconstructed image, and $\mathbf{y}^{k}$ are the respective undersampled k-space data.
@@ 3.2 Federated Generative Autoregressive Transformers
FedGAT is a model-agnostic FL technique that enables collaborative training of heterogeneous model architectures across sites by leveraging a decoupled two-tier strategy. During the first tier, using a server-client topology, a site-prompted global prior based on generative autoregressive transformers (GAT) is trained so as to capture the distribution of multi-site MR images (Fig. 1a). The FL server sends the trained global prior to individual sites at the end of the first tier. During the second tier, site-specific reconstruction models are locally trained with the aid of this global GAT prior (Fig. 1b). At each site, a reconstruction model of preferred architecture is pre-trained on the local dataset, and later fine-tuned on a hybrid dataset containing both the local dataset and synthetic datasets from other sites generated via the prior. The architecture of the site-prompted GAT prior, and image synthesis and two-tier training procedures for FedGAT are detailed in subsections below.
@@ 3.2.1 Architecture of the Site-Prompted GAT Prior
Figure 2: Architecture of the proposed site-prompted GAT prior. (a) The GAT prior embodies a variational autoencoder (VAE), whose encoder module maps an input MR image onto a set of discrete token maps $\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{S}$ across $S$ spatial scales, and whose decoder module recovers the input MR image back from the derived token maps. (b) The GAT prior also embodies a transformer module that builds an autoregressive prior on the multi-scale discrete token maps captured by the VAE. This module expresses generation of multi-scale discrete token maps as an autoregressive process across spatial scales, wherein a higher-scale token map $\mathbf{f}_{s}$ is predicted given the collection of token maps at earlier scales $\mathbf{f}_{<s}$:=assign\,:=\,:=$\{\mathbf{f}_{1},...,\mathbf{f}_{s-1}\}$. To preserve site-specific attributes in synthetic MR images, a site-prompt $\mathbf{sp}(k)$ is learned via a gated MLP (gMLP) network from a one-hot encoding of site index, and used to initialize the site token $\mathbf{st}$ in the autoregressive transformer module.
@@
Autoregressive Transformer Module: While the VAE encoder extracts discrete token maps residually across spatial scales, it does not explicitly examine the statistical relationships among these multi-scale maps [65]. To obtain a correlate for the joint distribution of $\mathbf{f}_{1},..,\mathbf{f}_{S}$, here we propose to build an autoregressive prior on multi-scale discrete token maps via the autoregressive transformer module. In particular, the transformer module aims to predict the sequence of tokens $\mathbf{f}_{s}$ at the $s$th scale, given as input the sequence of tokens aggregated across lower scales $\mathbf{f}_{<s}$:=assign\,:=\,:=$\{\mathbf{f}_{1},...,\mathbf{f}_{s-1}\}\in[V]^{T_{s-1}}$ ($T_{s-1}=\sum_{o=1}^{s-1}h_{o}w_{o}$ denoting the number of tokens at lower scales). To maintain spatial correspondence across consecutive scales, the token maps in $\mathbf{f}_{<s}$ are individually embedded after upsampling to the dimensions of the subsequent scale:
@@
where $\mathbf{e}_{o}\in\mathbb{R}^{h_{o+1}w_{o+1}\times d}$ and $\mathbf{e}_{<s}$:=assign\,:=\,:=$\{\mathbf{e}_{1},...,\mathbf{e}_{s-1}\}\in\mathbb{R}^{(T_{s}-1)\times d}$.
To enable the transformer module to synthesize feature maps aligned with site-specific data distributions, a learnable site-prompt $\mathbf{sp}(k)\in\mathbb{R}^{1\times d}$ is generated by projecting a one-hot encoding of site index ($\mathbb{1}_{K}(k)$) through a gated multi-layer perceptron (gMLP) network:
$\mathbf{g}$ $=\mathrm{GELU}(\mathrm{Lin}(\mathbb{1}_{K}(k))),$ (12)
$\mathbf{g}_{u},\mathbf{g}_{v}$ $=\mathrm{Split}(\mathbf{g}),$ (13)
$\mathbf{sp}(k)$ $=\mathrm{Lin}(\mathrm{Lin}(\mathrm{LN}(\mathbf{g}_{u}))\odot\mathbf{g}_{v}),$ (14)
where GELU is Gaussian error linear unit activation, $\mathbf{g}$ is the hidden activation, $\mathbf{g}_{u}$ and $\mathbf{g}_{v}$ are the channel-split components of $\mathbf{g}$, Lin denotes a linear layer, and LN denotes layer normalization. The site prompt is then used to initialize a site token as $\mathbf{st}$=\,=\,=$\mathbf{sp}(k)$. Afterwards, the site token is pooled with a learnable projection of the embedded token maps and additively combined with learnable position and scale encodings:
@@
To predict $\mathbf{f}_{s}$, the hidden representation $\mathbf{h}_{0}$ is projected through $L$ transformer blocks, each containing multi-head self-attention (MHSA) and MLP layers interleaved with normalization and residual connections [64]. Yet, unlike normalization layers in conventional transformers, here we employ adaptive layer normalization (AdaLN) in order to induce site-specific processing of hidden representations [70]. In particular, we propose to modulate the statistics of $\mathbf{h}$ via learnable functions of the site token $\mathbf{st}$: